# Minimal ring extensions of the integers exhibiting Kochen-Specker contextuality

### v3

@@ Abstract.
This paper is a contribution to the algebraic study of contextuality in quantum theory. As an algebraic analogue of Kochen and Specker’s no-hidden-variables result, we investigate rational subrings over which the partial ring of $d\times d$ symmetric matrices ($d\geq 3$) admits no morphism to a commutative ring, which we view as an “algebraic hidden state.” For $d=3$, the minimal such ring is shown to be $\mathbb{Z}[1/6]$, while for $d\geq 6$ the minimal subring is $\mathbb{Z}$ itself. The proofs rely on the construction of new sets of integer vectors in dimensions 3 and 6 that have no Kochen-Specker coloring.
@@ 1. Introduction
Quantum contextuality [7] is a foundational feature of quantum theory, which has grown to a widely studied phenomenon since the work of Bell [4] and Kochen-Specker [10]. Recall that the observables of a quantum-mechanical system are represented by elements of a noncommutative algebra of operators. For a pair of observables that do not commute, the uncertainty principle forbids us from simultaneously measuring the values of these observables with perfect certainty. Conversely, a pair of observables is called commeasurable if they commute, so that there is no such restriction on their simultaneous measurement. Broadly speaking, contextuality in quantum mechanics is the principle that the value of a measured observable is dependent on the commeasurable observables that happen to be probed by a measurement apparatus.
Contextuality has been studied through a number of different mathematical lenses. In recent years this has included sheaf theory [2, 3], Boolean algebra [5, 1], and topology [12, 11, 13]. In this paper we take inspiration directly from Kochen and Specker’s initial treatment of contextuality in terms of partial algebras, leading to a ring-theoretic and number-theoretic investigation as we explain below.
In the foundations of quantum physics, a hidden variable theory is an attempt to reduce the statistical uncertainty of the outcomes of quantum measurements to a statistical average over classical “hidden variables” that behave deterministically. Kochen and Specker’s investigation of contextuality centered on the proof that certain noncontextual hidden variable theories are impossible to construct. Their formulation of noncontextual hidden variable theories has a strongly algebraic flavor. We recall some terminology from [10] in order to state their result.
Let $K$ be a commutative ring. A partial $K$-algebra is a set $A$ equipped with a reflexive, symmetric binary relation $\odot\subseteq A\times A$ called commeasurability, along with partially defined operations $+,\,\cdot\ \colon\odot\to A$ of addition and multiplication as well as a globally defined scalar multiplication $K\times A\to A$, and distinguished elements $0,1\in A$, subject to the following axioms:
@@
Assuming (P1)–(P3), condition (P4) is equivalent to the statement that every set $S\subseteq A$ of pairwise commeasurable elements is contained in a set $C\subseteq A$ of pairwise commeasurable elements such that the restriction of all operations makes $C$ into a commutative $K$-algebra with additive and multiplicative identities 00 and $1$.
@@
where $\mathcal{O}$ is the partial $\mathbb{R}$-algebra of observables on the system, $\Omega$ is a set (viewed as a space of hidden states [10, p. 62] on which the hidden variables are defined), and $\mathbb{R}^{\Omega}$ is the algebra of all real-valued functions on $\Omega$. Note that evaluation at each element $x\in\Omega$ composes with $h$ to produce a morphism of partial $\mathbb{R}$-algebras $\psi_{x}\colon\mathcal{O}\to\mathbb{R}$, which assigns a definite value to every observable in a way that preserves algebraic relations between commeasurable observables. Conversely, if sufficiently many such hidden pure state morphisms $\psi\colon\mathcal{O}\to\mathbb{R}$ of partial algebras can be produced to separate distinct observables, we may take the product over these morphisms to construct a hidden variable theory $h\colon\mathcal{O}\to\mathbb{R}^{\Omega}$.
Kochen and Specker proved the following obstruction to any such hidden variable theory.
Kochen-Specker Theorem ([10]).
Let $d\geq 3$ be an integer. There is no morphism of partial $\mathbb{R}$-algebras from $\mathbb{M}_{d}(\mathbb{R})$ to any commutative $\mathbb{R}$-algebra. Thus there is no Kochen-Specker hidden variable theory for $\mathcal{O}=\mathbb{M}_{d}(\mathbb{C})_{\mathrm{sa}}$.
This was achieved by showing the non-existence of certain colorings on the partial Boolean algebra of projections, which in turn was deduced from the uncolorability of vectors that represent rank-1 projections. (The definition of these colorings and of partial Boolean algebras will be recalled in later sections of this paper.) Furthermore, they exhibited [10, Section 6] a hidden variable theory in the case of dimension 2 by constructing an injective morphism of partial $\mathbb{R}$-algebras $\mathbb{M}_{2}(\mathbb{C})_{\mathrm{sa}}\to\mathbb{R}^{\Omega}$ where $\Omega=S^{2}\subseteq\mathbb{R}^{3}$ is the unit sphere.
The formulation of the Kochen-Specker Theorem suggests a purely algebraic treatment of contextuality in the setting of noncommutative algebra via partial rings, explored in [18, 5]. It was shown in [18] (see also [19]) that Kochen and Specker’s original result provides an obstruction to the existence of functors that extend the prime spectrum from commutative to noncommutative rings, in the sense that such functors assign the empty set to the rings $\mathbb{M}_{d}(\mathbb{C})$ for $d\geq 3$. These two results were connected by the spectrum of partial prime ideals.
The problem of extending this result to matrix rings over any base naturally led to a more detailed study of partial algebras in [5]. By a partial ring we simply mean a partial $\mathbb{Z}$-algebra. In that paper, the obstruction for spectrum functors was reduced to $\mathbb{M}_{d}(\mathbb{Z})$. However, it was also shown that symmetric integer matrices have an algebraic analogue of a Kochen-Specker hidden state, taking values in a finite field, described as follows.
@@
Let $R$ be a partial ring. An algebraic hidden state (with values in $C$) is a morphism of partial rings $\psi\colon R\to C$ to a commutative ring $C$.
For a commutative ring $K$, we let $\mathbb{M}_{d}(K)_{\mathrm{sym}}$ denote the set of symmetric $d\times d$ matrices over $K$; this forms a partial $K$-algebra with commeasurability again being commutativity, and all operations restricted from those of the full matrix algebra. In the case $d=3$, it was proved in [5] that:
@@
  • there is no morphism of partial rings from $\mathbb{M}_{3}(\mathbb{Z}[1/30])_{\mathrm{sym}}$ to any commutative ring.
@@
The proof of this result depends on the existence of uncolorable sets of integer vectors. For this reason, Section 2 is devoted to the study of Kochen-Specker colorability of certain sets of integer vectors. The case of dimension $d=3$ relies on a set constructed in [6] along with a new set constructed in Theorem 3. The case of $d\geq 4$ makes use of uncolorable sets in dimension 4 as in any of the references [8, 9, 16]. The proof for $d\geq 6$ depends on a set whose uncolorability is verified by computational methods.
@@ Acknowledgments
This work would not have been possible without the previous efforts of Nikhil Dasgupta, Michael Pun, Samuel Swain, and Ira Hanzhao Li. We are grateful for their contributions.
@@ 2. Kochen-Specker colorings of integer vectors
Kochen and Specker’s no-hidden-variables theorem was proved by demonstrating the existence of a set of 117 three-dimensional vectors for which there is no KS coloring. Since then, a wide variety of KS uncolorable vector sets of varying dimensions have been found. The proof of Theorem 1 relies on the existence of Kochen-Specker uncolorable sets of integer vectors, i.e. vectors whose entries are all integers. Some of the best-known KS vector sets in fact consist of integer vectors. This includes the collections of vectors in dimension 4 with entries in $\{-1,0,1\}\subseteq\mathbb{Z}$, given by Peres [16], Kernaghan [9], and Cabello, Estebaranz, and García-Alcaine [8]. Similarly, the smallest known uncolorable vector set in dimension 3, due to Conway and Kochen [17, pp. 114, 197], consists of integer vectors on the cube with vertices (±$(\pm 2,\pm 2,\pm 2)$. Another uncolorable set of 37 integer vectors in dimension 3 was produced by Bub [6]. More recently, large numbers of integer KS vector sets were produced computationally by Pavičić [14] and Pavičić-Megill [15].
@@
For instance, if $S_{1}$ denotes any of the sets of 4-dimensional integer vectors in [8, 9, 16], then satisfies $q(S_{1})=\{1,2,4\}$ and thus $N(S_{1})=4$. Similarly, for the set $S_{2}\subseteq\mathbb{Z}^{3}$ constructed in [6], we have $q(S_{2})=\{1,2,3,5,6,30\}$ so that $N(S_{2})=30$. The Conway-Kochen set $S_{3}\subseteq\mathbb{Z}^{3}$ is traditionally represented [17, p. 114] by vectors on the cube with vertices (±$(\pm 2,\pm 2,\pm 2)$, so that it satisfies $q(S_{3})=\{4,5,6,8,9,12\}$ and $N(S_{3})=360$. However, some of these vectors such as $(2,2,2)=2\cdot(1,1,1)$ or $(2,0,0)=2\cdot(1,0,0)$ can be replaced by a smaller collinear integer vector to obtain an equivalently uncolorable set $S_{4}\subseteq\mathbb{Z}^{3}$ that satisfies $q(\mathcal{S}_{4})=\{1,2,3,5,6,9\}$ and $N(S_{4})=90$.
@@
$\mathcal{S}_{d}(N)$ =$=\{v\in\mathbb{Z}^{d}:q(v)\mbox{ is a unit in }\mathbb{Z}[1/N]\}$
=$=\{v\in\mathbb{Z}^{d}:q(v)\mbox{ divides a power of }N\}.$
@@
$\mathcal{S}_{d}(N)\mbox{ colorable}$ $\implies\mathcal{S}_{d}(M)\mbox{ colorable,}$,
$\mathcal{S}_{d}(M)\mbox{ uncolorable}$ $\implies\mathcal{S}_{d}(N)\mbox{ uncolorable.}$.
@@ 2.1. Uncolorability in dimension 3
We first focus on the classical case of dimension $d=3$. It was shown in [5, Theorem 3.4] that if $N$ is not divisible by 2 or not divisible by 3, then $\mathcal{S}_{3}(N)$ has a KS coloring. As an immediate consequence,
@@
In light of the above facts, one might naturally guess that $\mathcal{S}_{3}(6)$ has no Kochen-Specker coloring. However, to date we have not managed to find a proof or refutation of this statement. Lacking such an answer, we wondered whether there exists an integer $N$, divisble by $6$ but not $5$, such that $\mathcal{S}_{3}(N)$ is KS uncolorable. In this subsection we exhibit just such an example, for the specific value
@@
This will be proved by constructing an uncolorable subset $Q\subseteq\mathcal{S}_{3}(462)$, which we now describe. When considering Kochen-Specker colorings of sets of vectors, each vector is intended to represent the rank-one projection onto the line through that vector. For this reason it suffices to consider sets of noncollinear vectors. When restricting attention to integer vectors, this means that we may consider only those vectors whose entries have greatest common divisor equal to 1. We will call such integer vectors primitive.
@@
For instance, the vectors $(1,0,0)$, $(0,1,-1)$, $(1,1,1)$, and $(1,-1,1)$ are well-signed while $(-1,0,0)$, $(0,-1,1)$, $(-1,-1,-1)$, and $(-1,1,-1)$ are not. It is clear that for each $v\in\mathbb{Z}^{3}\setminus\{0\}$, exactly one of $v$ or $-v$ is well-signed. It follows that if a set $S\subseteq\mathbb{Z}^{3}$ consists of primitive, well-signed vectors, then the vectors in $S$ must be noncollinear.
@@
This is a set of $85$ integer vectors that are primitive, well-signed, and thus noncollinear. This set and its uncolorability were discovered using Mathematica software. [^1] The code is publicly available at https://github.com/manny-reyes/Kochen_Specker_Colorability. However, we are able to produce a human-readable proof of this fact as follows.
@@
the vector $(1,0,0)$ is colored $1$ while the other two are colored 00. It follows that in the triple
@@
either the second or third vector must be colored 1. Because multiplication by the diagonal reflection matrix $R_{z}=\operatorname{diag}(1,1,-1)$ leaves the coordinate axes invariant while interchanging the vectors $(1,0,\pm 1)$, we may again assume without loss of generality that $(1,0,1)$ is colored 0 and $(1,0,-1)$ is colored 1. Finally, in the orthogonal triple
@@
either the second or third vector must be colored 1. Since multiplication by the reflection matrix $R_{y}=\operatorname{diag}(1,-1,1)$ again leaves the coordinate axes and the lines through $(1,0,\pm 1)$ invariant while interchanging the vectors $(1,\pm 1,0)$, we may assume without loss of generality that $(1,1,0)$ is assigned 0 and $(1,-1,0)$ is 1.
@@
[…]
@@
[…]
But this contradicts the previous deduction that $(-3,8,2)$ must be colored 0, completing the proof. ∎

### v4

@@ Abstract.
This paper is a contribution to the algebraic study of contextuality in quantum theory. As an algebraic analogue of Kochen and Specker’s no-hidden-variables result, we investigate rational subrings over which the partial ring of $d\times d$ symmetric matrices ($d\geq 3$) admits no morphism to a nonzero commutative ring, which we view as an “algebraic hidden state.” For $d=3$, the minimal such ring is shown to be $\mathbb{Z}[1/6]$, while for $d\geq 6$ the minimal subring is $\mathbb{Z}$ itself. The proofs rely on the construction of new sets of integer vectors in dimensions 3 and 6 that have no Kochen-Specker coloring.
@@ 1. Introduction
Quantum contextuality [BCGKL] is a foundational feature of quantum theory, which has grown to a widely studied phenomenon since the work of Bell [Bell] and Kochen-Specker [KochenSpecker]. Recall that the observables of a quantum-mechanical system are represented by elements of a noncommutative algebra of operators. For a pair of observables that do not commute, the uncertainty principle forbids us from simultaneously measuring the values of these observables with perfect certainty. Conversely, a pair of observables is called commeasurable if they commute, so that there is no such restriction on their simultaneous measurement. Broadly speaking, contextuality in quantum mechanics is the principle that the value of a measured observable is dependent on the commeasurable observables that happen to be probed by a measurement apparatus.
Contextuality has been studied through a number of different mathematical lenses. In recent years this has included sheaf theory [AbramskyBrandenburger, AMB], Boolean algebra [BMR, AbramskyBarbosa], and topology [ORBR, OkayRaussendorf, OkaySheinbaum]. In this paper we take inspiration directly from Kochen and Specker’s initial treatment of contextuality in terms of partial algebras, leading to a ring-theoretic and number-theoretic investigation as we explain below.
In the foundations of quantum physics, a hidden variable theory is an attempt to reduce the uncertainty of the outcomes of quantum measurements to a statistical average over classical “hidden variables” that behave deterministically. Kochen and Specker’s investigation of contextuality centered on the proof that certain noncontextual hidden variable theories are impossible to construct. Their formulation of noncontextual hidden variable theories has a strongly algebraic flavor. We recall some terminology from [KochenSpecker] in order to state their result.
Let $K$ be a commutative ring. A partial $K$-algebra is a set $A$ equipped with a reflexive, symmetric binary relation $\odot\subseteq A\times A$ called commeasurability, along with partially defined operations $\cdot\ \colon\odot\to A$ of addition and multiplication as well as a globally defined scalar multiplication $K\times A\to A$, and distinguished elements $0,1\in A$, subject to the following axioms:
@@
Assuming (P1)–(P3), condition (P4) is equivalent to the statement that every set $S\subseteq A$ of pairwise commeasurable elements is contained in a set $C\subseteq A$ of pairwise commeasurable elements such that the restriction of all operations makes $C$ into a commutative $K$-algebra with additive and multiplicative identities 0 and $1$.
@@
where $\mathcal{O}$ is the partial $\mathbb{R}$-algebra of observables on the system, $\Omega$ is a set (viewed as a space of hidden states [KochenSpecker, p. 62] on which the hidden variables are defined), and $\mathbb{R}^{\Omega}$ is the algebra of all real-valued functions on $\Omega$. Note that evaluation at each element $x\in\Omega$ composes with $h$ to produce a morphism of partial $\mathbb{R}$-algebras $\psi_{x}\colon\mathcal{O}\to\mathbb{R}$, which assigns a definite value to every observable in a way that preserves algebraic relations between commeasurable observables. Conversely, if sufficiently many such hidden pure state morphisms $\psi\colon\mathcal{O}\to\mathbb{R}$ of partial algebras can be produced to separate distinct observables, we may take the product over these morphisms to construct a hidden variable theory $h\colon\mathcal{O}\to\mathbb{R}^{\Omega}$.
Kochen and Specker proved [KochenSpecker, pp. 66 & 70] the following obstruction to any such hidden variable theory.
Kochen-Specker Theorem ([KochenSpecker]).
Let $d\geq 3$ be an integer. There is no morphism of partial $\mathbb{R}$-algebras from $\mathbb{M}_{d}(\mathbb{R})_{\mathrm{sym}}$ to any nonzero commutative $\mathbb{R}$-algebra. Thus there is no Kochen-Specker hidden variable theory for $\mathcal{O}=\mathbb{M}_{d}(\mathbb{C})_{\mathrm{sa}}$.
This was achieved by showing the non-existence of certain colorings on the partial Boolean algebra of projections, which in turn was deduced from the uncolorability of vectors that represent rank-1 projections. (The definition of these colorings and of partial Boolean algebras will be recalled in later sections of this paper.) Furthermore, they exhibited [KochenSpecker, Section 6] a hidden variable theory in the case of dimension 2 by constructing an injective morphism of partial $\mathbb{R}$-algebras $\mathbb{M}_{2}(\mathbb{C})_{\mathrm{sa}}\to\mathbb{R}^{\Omega}$ where $\Omega=S^{2}\subseteq\mathbb{R}^{3}$ is the unit sphere.
The formulation of the Kochen-Specker Theorem suggests a purely algebraic treatment of contextuality in the setting of noncommutative algebra via partial rings, explored in [Reyes, BMR]. It was shown in [Reyes] (see also [Reyes:tour]) that Kochen and Specker’s original result provides an obstruction to the existence of functors that extend the prime spectrum from commutative to noncommutative rings, in the sense that such functors assign the empty set to the rings $\mathbb{M}_{d}(\mathbb{C})$ for $d\geq 3$. These two results were connected by the spectrum of partial prime ideals.
The problem of extending this result to matrix rings over any base naturally led to a more detailed study of partial algebras in [BMR]. By a partial ring we simply mean a partial $\mathbb{Z}$-algebra. In that paper, the obstruction for spectrum functors was reduced to $\mathbb{M}_{d}(\mathbb{Z})$. However, it was also shown that symmetric integer matrices have an algebraic analogue of a Kochen-Specker hidden state, taking values in a finite field, described as follows.
@@
Let $R$ be a partial ring. An algebraic hidden state (with values in $C$) for $R$ is a morphism of partial rings $\psi\colon R\to C$ to a nonzero commutative ring $C$.
For a commutative ring $K$, we let $\mathbb{M}_{d}(K)_{\mathrm{sym}}$ denote the set of symmetric $d\times d$ matrices over $K$; this forms a partial $K$-algebra with commeasurability again being commutativity, and all operations restricted from those of the full matrix algebra. In the case $d=3$, it was proved in [BMR] that:
@@
  • there is no morphism of partial rings from $\mathbb{M}_{3}(\mathbb{Z}[1/30])_{\mathrm{sym}}$ to any nonzero commutative ring.
@@
The proof of this result depends on the existence of uncolorable sets of integer vectors. For this reason, Section 2 is devoted to the study of Kochen-Specker colorability of certain sets of integer vectors. The case of dimension $d=3$ relies on a set constructed in [Bub:Schutte] along with a new set constructed in Theorem 3. The case of $d\geq 4$ makes use of uncolorable sets in dimension 4 as in any of the references [CEG, Kernaghan, Peres:KS]. The proof for $d\geq 6$ depends on a set whose uncolorability is verified by computational methods.
@@ Acknowledgments
This work would not have been possible without the previous efforts of Nikhil Dasgupta, Michael Pun, Samuel Swain, and Ira Hanzhao Li, and we are grateful for their contributions. We also thank the anonymous referee for some helpful corrections.
@@ 2. Kochen-Specker colorings of integer vectors
Kochen and Specker’s no-hidden-variables theorem was proved by demonstrating the existence of a set of 117 three-dimensional vectors for which there is no KS coloring. Since then, a wide variety of KS uncolorable vector sets of varying dimensions have been found. The proof of Theorem 1 relies on the existence of Kochen-Specker uncolorable sets of integer vectors, i.e. vectors whose entries are all integers. Some of the best-known KS vector sets in fact consist of integer vectors. This includes the collections of vectors in dimension 4 with entries in $\{-1,0,1\}\subseteq\mathbb{Z}$, given by Peres [Peres:KS], Kernaghan [Kernaghan], and Cabello, Estebaranz, and García-Alcaine [CEG]. Similarly, the smallest known uncolorable vector set in dimension 3, due to Conway and Kochen [Peres:quantum, pp. 114, 197], consists of integer vectors on the cube with vertices $(\pm 2,\pm 2,\pm 2)$. Another uncolorable set of 37 integer vectors in dimension 3 was produced by Bub [Bub:Schutte]. More recently, large numbers of integer KS vector sets were produced computationally by Pavičić [Pavicic:exhaustive] and Pavičić-Megill [PavicicMegill:automated].
@@
For instance, if $S_{1}$ denotes any of the sets of 4-dimensional integer vectors in [CEG, Kernaghan, Peres:KS], then satisfies $q(S_{1})=\{1,2,4\}$ and thus $N(S_{1})=4$. Similarly, for the set $S_{2}\subseteq\mathbb{Z}^{3}$ constructed in [Bub:Schutte], we have $q(S_{2})=\{1,2,3,5,6,30\}$ so that $N(S_{2})=30$. The Conway-Kochen set $S_{3}\subseteq\mathbb{Z}^{3}$ is traditionally represented [Peres:quantum, p. 114] by vectors on the cube with vertices $(\pm 2,\pm 2,\pm 2)$, so that it satisfies $q(S_{3})=\{4,5,6,8,9,12\}$ and $N(S_{3})=360$. However, some of these vectors such as $(2,2,2)=2\cdot(1,1,1)$ or $(2,0,0)=2\cdot(1,0,0)$ can be replaced by a smaller collinear integer vector to obtain an equivalently uncolorable set $S_{4}\subseteq\mathbb{Z}^{3}$ that satisfies $q(\mathcal{S}_{4})=\{1,2,3,5,6,9\}$ and $N(S_{4})=90$.
@@
$\mathcal{S}_{d}(N)$ $=\{v\in\mathbb{Z}^{d}:q(v)\mbox{ is a unit in }\mathbb{Z}[1/N]\}$
$=\{v\in\mathbb{Z}^{d}:q(v)\mbox{ divides a power of }N\}.$
@@
$\mathcal{S}_{d}(N)\mbox{ colorable}$ $\implies\mathcal{S}_{d}(M)\mbox{ colorable,}$
$\mathcal{S}_{d}(M)\mbox{ uncolorable}$ $\implies\mathcal{S}_{d}(N)\mbox{ uncolorable.}$
@@ 2.1. Uncolorability in dimension 3
We first focus on the classical case of dimension $d=3$. It was shown in [BMR, Theorem 3.4] that if $N$ is not divisible by 2 or not divisible by 3, then $\mathcal{S}_{3}(N)$ has a KS coloring. As an immediate consequence,
@@
In light of the above facts, one might naturally guess that $\mathcal{S}_{3}(6)$ has no Kochen-Specker coloring. However, to date we have not managed to find a proof or refutation of this statement. Lacking such an answer, we wondered whether there exists an integer $N$, divisible by $6$ but not $5$, such that $\mathcal{S}_{3}(N)$ is KS uncolorable. In this subsection we exhibit just such an example, for the specific value
@@
This will be proved by constructing an uncolorable subset $Q\subseteq\mathcal{S}_{3}(462)$, which we now describe. When considering Kochen-Specker colorings of sets of vectors, each vector is intended to represent the rank-one projection onto the line through that vector. For this reason it suffices to consider sets of pairwise noncollinear vectors. When restricting attention to integer vectors, this means that we may consider only those vectors whose entries have greatest common divisor equal to 1. We will call such integer vectors primitive.
@@
For instance, the vectors $(1,0,0)$, (0,1,1)(0,1,-1), $(1,1,1)$, and (1,1,1)(1,-1,1) are well-signed while (1,0,0)(-1,0,0), (0,1,1)(0,-1,1), (1,1,1)(-1,-1,-1), and (1,1,1)(-1,1,-1) are not. It is clear that for each $v\in\mathbb{Z}^{3}\setminus\{0\}$, exactly one of $v$ or v-v is well-signed. It follows that if a set $S\subseteq\mathbb{Z}^{3}$ consists of primitive, well-signed vectors, then the vectors in $S$ must be noncollinear.
@@
This is a set of $85$ integer vectors that are primitive, well-signed, and thus noncollinear. This set and its uncolorability were discovered using Mathematica software. [^1] The code is publicly available at https://github.com/manny-reyes/Kochen_Specker_Colorability/tree/main/462_vector_set. However, we are able to produce a human-readable proof of this fact as follows.
@@
the vector $(1,0,0)$ is colored $1$ while the other two are colored 0. It follows that in the triple
@@
either the second or third vector must be colored 1. Because multiplication by the diagonal reflection matrix $R_{z}=\operatorname{diag}(1,1,-1)$ leaves the coordinate axes invariant while interchanging the vectors $(1,0,\pm 1)$, we may again assume without loss of generality that $(1,0,1)$ is colored 0 and (1,0,1)(1,0,-1) is colored 1. Finally, in the orthogonal triple
@@
either the second or third vector must be colored 1. Since multiplication by the reflection matrix $R_{y}=\operatorname{diag}(1,-1,1)$ again leaves the coordinate axes and the lines through $(1,0,\pm 1)$ invariant while interchanging the vectors $(1,\pm 1,0)$, we may assume without loss of generality that $(1,1,0)$ is assigned 0 and (1,1,0)(1,-1,0) is 1.
@@
$\{(1,0,-1),\ (1,-1,1)\}\mbox{ orthogonal}$ $\implies(1,-1,1)\mapsto 0,$
$\{(0,1,1),\ (1,-1,1),\ (2,1,-1)\}\mbox{ orthogonal}$ $\implies(2,1,-1)\mapsto 1,$
$\{(2,1,-1),(-3,8,2)\}\mbox{ orthogonal}$ $\implies(-3,8,2)\mapsto 0.$
@@
$\{(1,-1,0),\ (1,1,-1)\}\mbox{ orthogonal}$ $\implies(1,1,-1)\mapsto 0,$
$\{(1,0,1),\ (1,1,-1),\ (-1,2,1)\}\mbox{ orthogonal}$ $\implies(-1,2,1)\mapsto 1,$
$\{(-1,2,1),\ (4,1,2)\}\mbox{ orthogonal}$ $\implies(4,1,2)\mapsto 0,$
$\{(1,-1,0),\ (2,2,-5)\}\mbox{ orthogonal}$ $\implies(2,2,-5)\mapsto 0,$
$\{(4,1,2),\ (2,2,-5),\ (-3,8,2)\}\mbox{ orthogonal}$ $\implies(-3,8,2)\mapsto 1.$
But this contradicts the previous deduction that (3,8,2)(-3,8,2) must be colored 0, completing the proof. ∎